According to a candy company, packages of a certain candy contain 17% orange candies. Suppose we examine 100 random candies.
A. What is the standard error? Round to three decimal places
B. We expect (blank) % orange candies, give or take (blank) round to one decimal place
thank you!

Answer by stanbon(75887)

According to a candy company, packages of a certain candy contain 17% orange candies. Suppose we examine 100 random candies.
A. What is the standard error? Round to three decimal places
standard error = sqrt[pq/n] = sqrt[0.17*0.83/100] = 0.0376
---------------------------------------------------
B. We expect (blank) % orange candies, give or take (blank) round to one decimal place
mean = np = 100*0.17 = 17
----------------
Cheers,
Stan H.
-----------------------